OpenStudy (anonymous):

find the roots of 3x^2=-48?? Please :)

OpenStudy (anonymous):

3x+48=0

OpenStudy (anonymous):

\[3x^2=-48 \rightarrow 3 x^2+48 = 0\]

OpenStudy (anonymous):

solutions:\[x = -4 i\] solutions: \[x = 4 i\]

OpenStudy (anonymous):

how did you get x=-4i and x=4i?

OpenStudy (anonymous):

\[3 x^2 = -48\] \[3 x^2+48 = 0\]

OpenStudy (anonymous):



OpenStudy (anonymous):